Sylvia and Jadon now want to work a problem. Imagine a puck of mass 0.5 kg moving as in the simulation. Suppose that the tension in the string is 2.0 N, and that the radius of its circular path is 0.9 m. What will Jadon and Sylvia find for the tangential speed of the puck?



__ m/s


1
Expert's answer
2021-10-11T16:23:32-0400

"F=\frac{mv^2}{r}\\v=\sqrt\frac{Fr}{m}"

"v=\sqrt\frac{2\times0.9}{0.5}"

v=1.89m/sec